### v1

# RefineVAD: Semantic-Guided Feature Recalibration for Weakly Supervised Video Anomaly Detection

Abstract
Weakly-Supervised Video Anomaly Detection aims to identify anomalous events using only video-level labels, balancing annotation efficiency with practical applicability. However, existing methods often oversimplify the anomaly space by treating all abnormal events as a single category, overlooking the diverse semantic and temporal characteristics intrinsic to real-world anomalies. Inspired by how humans perceive anomalies, by jointly interpreting temporal motion patterns and semantic structures underlying different anomaly types, we propose RefineVAD, a novel framework that mimics this dual-process reasoning. Our framework integrates two core modules. The first, Motion-aware Temporal Attention and Recalibration (MoTAR), estimates motion salience and dynamically adjusts temporal focus via shift-based attention and global Transformer-based modeling. The second, Category-Oriented Refinement (CORE), injects soft anomaly category priors into the representation space by aligning segment-level features with learnable category prototypes through cross-attention. By jointly leveraging temporal dynamics and semantic structure, explicitly models both “how” motion evolves and “what” semantic category it resembles. Extensive experiments on WVAD benchmark validate the effectiveness of RefineVAD and highlight the importance of integrating semantic context to guide feature refinement toward anomaly-relevant patterns.
Figure 1: Comparison of WVAD paradigms. (a) Coarse-grained model predicts anomaly scores per snippet. (b) Fine-grained method introduces auxiliary video-level category classification, but these labels are not utilized during anomaly scoring. (c) RefineVAD (Ours): category-aware soft classification guides feature enhancement with learnable prototypes.
Codehttps://github.com/VisualScienceLab-KHU/RefineVAD
Introduction
Video Anomaly Detection (VAD) aims to identify unexpected or abnormal events within video streams and has become increasingly important in applications such as surveillance, public safety, and industrial monitoring. While fully supervised methods achieve precise anomaly localization, they require costly frame-level annotations, limiting their scalability in real-world scenarios. Unsupervised approaches, which learn solely from normal patterns, reduce labeling costs but often suffer from poor generalization and high false-positive rates. To balance annotation efficiency with detection performance, Weakly Supervised Video Anomaly Detection (WVAD) has emerged as a compelling alternative. WVAD methods rely exclusively on video-level labels—indicating whether a video contains any anomaly—and typically adopt a Multiple Instance Learning (MIL) paradigm, wherein each video is treated as a bag of segments, and at least one segment in an anomalous video is assumed to contain abnormal content.
However, recognizing abnormal events in videos requires more than coarse video-level classification. Human perception leverages two complementary dimensions when detecting anomalies: (1) contextual motion dynamics that evolve over time—such as abrupt or irregular movements—and (2) prior knowledge of what type of anomaly is occurring, allowing us to interpret visual patterns semantically. This motivates a hierarchical reasoning framework that first attends to how motion unfolds and then reasons about what kind of abnormal event it resembles. Such reasoning enables more robust and interpretable anomaly detection by aligning perceptual cues with semantic understanding.
Despite recent advancements, existing WVAD methods exhibit two critical limitations that impede this hierarchical reasoning. First, temporal modeling is often shallow or inflexible, relying on fixed pooling operations or simple aggregation schemes that fail to adapt to the diverse motion characteristics of real-world anomalies. Since many anomalies are defined by dynamic, non-uniform, or context-specific motion patterns, this temporal rigidity significantly hampers localization accuracy. Second, and more fundamentally, most WVAD frameworks treat all abnormal events as a single generic class, ignoring the semantic diversity across anomaly types. While the original VAD task framed anomaly detection as a binary classification problem—determining whether each snippet is normal or abnormal, in practice, anomalies exhibit highly distinctive visual and motion patterns. For instance, fighting involves abrupt, bidirectional motion, while explosions are characterized by sudden flashes and spatial bursts. Ignoring such distinctions limits the model’s ability to capture discriminative features, thereby reducing both interpretability and localization performance. Although some recent works have attempted to incorporate anomaly categories, these are typically used only for auxiliary video-level classification and not integrated into the anomaly scoring or representation learning processes. Consequently, semantic distinctions remain unexploited during temporal localization, and the model fails to learn category-aware features crucial for precise detection.
To address these challenges, we propose RefineVAD, a novel framework that introduces motion-aware temporal recalibration and category-conditioned refinement into the anomaly detection process. RefineVAD integrates two key components—MoTAR and CORE—to explicitly models both “how” motion evolves and “what” semantic category it resembles, enabling structured anomaly reasoning under weak supervision. The Motion-aware Temporal Attention and Recalibration (MoTAR) module captures motion-intensity-aware relationships between segments by dynamically shifting feature channels in proportion to estimated motion salience. It further employs a lightweight Transformer-based global attention mechanism to model long-range dependencies, allowing the model to focus on critical temporal regions without relying on rigid windowing or fixed aggregation.
The Category-Oriented Refinement (CORE) module incorporates semantic priors by selectively injecting soft category prototypes into the representation space. It first predicts a video-level anomaly category distribution via soft classification, which is then used to compute a weighted combination of learnable category embeddings. Each embedding serves as a prototype encoding the characteristic patterns of a specific anomaly type. These category-aware combined embeddings are injected into the snippet-level features via cross-attention, guiding the model to focus on temporally and semantically meaningful regions during scoring. By combining these two components, RefineVAD effectively disentangles and recomposes multi-modal evidence to enhance anomaly localization. It captures the temporal evolution of motion and aligns it with category-relevant semantics, thereby improving both detection precision and interpretability.
Our contributions are summarized as follows:
  • We propose RefineVAD, a novel WVAD framework that jointly models temporal and semantic contexts for anomaly localization under weak supervision.
  • We design MoTAR, a motion-aware temporal attention module that adaptively recalibrates temporal feature flows based on motion salience.
  • We introduce CORE, a category-oriented refinement module that injects soft category priors into snippet-level representations via attention-guided semantic alignment.
  • We demonstrate that RefineVAD significantly outperforms previous state-of-the-art methods on multiple benchmark datasets through comprehensive ablation studies.
Related Works
Weakly Supervised Video Anomaly Detection.
VAD aims to identify events that deviate from typical or expected normal patterns in video streams. Traditional fully supervised learning methods detect anomalies using frame-level annotations, which are expensive and difficult to apply at scale  (zhu2021video; majhi2021weakly; majhi2021dam). On the other hand, Unsupervised VAD (UVAD) methods trained only on normal data struggle to robustly characterize normality and lead to reconstruction and prediction errors  (hasan2016learning; liu2018future; wu2019deep; cai2021appearance). To bridge this gap, WVAD has emerged and significantly reduce the annotation cost since it only requires video-level labels to indicate whether anomalous signs are present in the video  (wu2020not; zaheer2020claws; tian2021weakly; chen2023tevad; cho2023look; wan2020weakly). Most WVAD methods are based on the Multiple-Instance Learning (MIL) framework to utilize anomalous video labels without frame annotations. A foundational approach applies a ranking loss to enforce higher scores for anomalous snippets compared to normal ones (sultani2018real). Subsequent works build upon this paradigm by introducing techniques such as two-stage pseudo-labeling or surrogate optimization (feng2021mist). These MIL methods typically follow a pipeline that extracts pre-trained features (e.g., C3D (tran2015learning)/I3D (carreira2017quo)) and then trains a snippet classifier.
Figure 2: The overall architecture of proposed RefineVAD. Visual and textual features are extracted from each segment and fused before being processed by MoTAR, which adaptively recalibrates temporal features based on motion salience. The resulting features are then refined by CORE, which injects soft category priors via cross-attention with learnable prototypes, enabling category-aware anomaly localization.
Category and Semantic-Guided VAD.
More recent approaches shift towards leveraging semantic priors and category awareness to enhance weakly supervised video anomaly detection. VadCLIP (wu2024vadclip) applies frozen CLIP(radford2021clip) to WSVAD via a dual-branch design: one branch handles binary classification, while the other aligns frame features with text label embeddings. This enables both coarse- and fine-grained anomaly detection. TPWNG (yang2024text) leverages CLIP with learnable prompts and event categories as text inputs, generating snippet-level pseudo-labels and training a video classifier in a self-learning loop with temporal context modeling. PEMIL (chen2024prompt) formulates WSVAD as prompt learning, injecting semantic priors via anomaly and normal context prompts constructed from class label words. While these methods effectively exploit CLIP’s vision-language capability, they rely on discrete category labels or handcrafted prompts, which may limit adaptability to ambiguous or overlapping anomalies. Moreover, they often overlook inter-category dependencies or the channel-wise nature of visual representations.
Proposed Methods
To address the limitations of existing WVAD methods—namely, their static temporal modeling and disregard for semantic category structures—we propose RefineVAD, a modular framework that introduces contextual awareness at two complementary levels: (1) motion-aware temporal recalibration, and (2) category-conditioned semantic refinement. Rather than relying solely on visual embeddings or treating all anomalies as a single generic class, RefineVAD leverages both structured visual and textual information, dynamically models motion salience over time, and integrates category-aware priors to enhance anomaly representation and scoring.
As illustrated in Figure 2, RefineVAD follows the MIL setup by dividing each input video—annotated with a video-level anomaly label—into $T$ fixed-length segments. Each segment is independently encoded using pretrained visual and textual encoders to extract modality-specific features. These features are then concatenated to form a joint multi-modal representation for each segment, capturing both appearance and semantic context. This joint representation is first processed by the MoTAR (Motion-aware Temporal Attention and Recalibration) module. MoTAR dynamically adjusts feature channel shift ratios based on segment motion intensity to achieve adaptive temporal recalibration based on the degree of motion irregularity. In parallel, it captures long-range temporal dependencies using a lightweight Transformer-based attention mechanism (vaswani2017attention), allowing flexible and adaptive modeling of without rigid pooling or fixed windowing.
Next, the motion-contextual features are refined by the CORE (Category-Oriented Refinement) module, which incorporates semantic priors into the representation space. CORE first performs a soft classification at the video-level to estimate the probability distribution over anomaly categories. These probabilities are used to compute a weighted combination of learnable category prototypes, each encoding representative patterns of a specific anomaly type. The resulting soft category embedding is then injected into all segment-level features via cross-attention, enabling the model to emphasize segments that are not only temporally salient but also semantically consistent with the predicted anomaly type.
By jointly modeling motion intensity and category relevance, RefineVAD enables the disentanglement and recomposition of multi-modal cues to enhance anomaly localization. The final category-aware, temporally contextualized features are passed through a lightweight classifier to compute segment-level anomaly scores, which are subsequently aggregated to produce the video-level anomaly prediction.
Motion-aware Temporal Attention Recalibration
Figure 3: Overview of MoTAR. Motion variance guides adaptive channel shifting, enabling dynamic temporal feature aggregation. A Global Transformer captures long-range dependencies, improving sensitivity to diverse motions.
Anomalous events in videos often exhibit irregular motion dynamics—some are abrupt, while others remain deceptively static. To accommodate this variability, we propose MoTAR, a lightweight and effective temporal feature modeling module.
MoTAR is built upon the Temporal Shift Module (TSM) (lin2019tsm), which introduces temporal interactions by shifting a portion of feature channels forward and backward in time. While TSM is widely used in video understanding due to its efficiency and parameter free design, it applies a fixed shift ratio across all frames, limiting its adaptability to frames with varying motion intensities.
To address this, MoTAR first estimates the motion intensity per-frame using feature differences. As shown in Figure 3, given an input from the vision and text encoder $\mathbf{X}=[\mathbf{x}_{1},\dots,\mathbf{x}_{T}]\in\mathbb{R}^{T\times D}$, where $T$ is the number of sequence and $\mathbf{D}$ is the feature dimension, we compute $\Delta_{t}=\mathbf{x}_{t}-\mathbf{x}_{t-1}$ and obtain $\mathbf{v_{t}}=\mathrm{Var}(\Delta_{t})$ where $\mathbf{v_{t}}$ reflects local motion intensity. A higher variance indicates more significant motion, suggesting the need for broader temporal context aggregation. This variance-based measure is representation-agnostic, noise-robust, and enables adaptive temporal shifting with minimal computational overhead.
The variance vector is processed by a lightweight MLP-based shift predictor to obtain a shift ratio $r_{t}\in[0,1]$:
$r_{t}=\sigma(W_{3}\cdot\phi(W_{2}\cdot\phi(W_{1}v_{t})))$ (1)
where $\phi$ denotes the GELU activation and $\sigma$ is the sigmoid function.The number of channels to shift across features is given by:
$s_{t}=\left\lfloor r_{t}\cdot\frac{D}{K}\right\rfloor$ (2)
where $K$ is a predefined folding factor. The temporally shifted output of the $t$-th segment is then constructed as:
$\mathbf{y}_{t}=\begin{bmatrix}\mathbf{x}_{t-1}^{(1:s_{t})}&\mathbf{x}_{t+1}^{(s_{t}:2s_{t})}&\mathbf{x}_{t}^{(2s_{t}:D)}\end{bmatrix}$ (3)
This dynamic construction allows the model to allocate more shifted channels to high-motion frames, while preserving static content in low-variance frames.
We further apply a lightweight Transformer encoder over the shifted sequence to incorporate long-range temporal dependencies, producing temporal-contextual features.:
$\mathbf{X}^{\text{tc}}=\mathrm{Transformer}([\mathbf{y}_{1},\dots,\mathbf{y}_{T}])$ (4)
In summary, MoTAR extracts temporal-contextual feature alignments through adaptive shifting and capturing global temporal context. Its parameter-free shifting ensures computational efficiency, making it suitable for real-time or large-scale video anomaly detection applications.
Detailed numerical results and a comprehensive analysis of the computational efficiency are provided in Appendix F.
Category-Oriented Refinement
To enrich temporal contextual feature with semantic information, we introduce CORE module, a key component of the RefineVAD framework. CORE systematically injects category‑specific semantics into segment-level features, thereby guiding the model to better focus on distinctive aspects of each abnormal event through semantic reasoning. The module consists of two main stages. (1) Soft Category Classification: We first distinguish the anomaly category associated with a given video sample by applying a soft classification mechanism, which estimates the probability of the sample belonging to each predefined anomaly class. (2) Category Prototype Injection: These probabilities are then used as weights to softly combine category-specific prototype embeddings, resulting in a unified semantic vector that captures both dominant and complementary category traits. This vector is further injected into individual temporal features via a cross‑attention mechanism, producing category‑aware refined features that dynamically align each segment with the most relevant category semantics. By enriching the representation with this semantically guided context, the model achieves improved discriminability in the final stage of anomaly scoring.
Soft Category Classification. Let $\mathbf{X}^{\text{tc}}=[\mathbf{x}_{1},\dots,\mathbf{x}_{T}]$ be the temporal-contextual feature sequence generated by the MoTAR module, where each $\mathbf{x}_{t}^{\text{tc}}\in\mathbb{R}^{D}$ corresponds to the contextualized feature of the $t$-th segment. Each segment-level feature $\mathbf{x}_{t}^{\text{tc}}$ is processed by relatively simple fully connected (FC) layers consisting of dropout and batch normalization layers for scoring. These roughly estimated anomaly scores are converted into normalized weights which are used to aggregate the segment features along the temporal dimension. Since anomalous regions are relatively sparse within a video, it is more advantageous to focus on high-confidence segments rather than the entire video when distinguishing between categories.
The pooled video feature is passed through a soft category classifier, yielding logits that we reshape into a matrix $\mathbf{z}\in\mathbb{R}^{\mathbf{C}\times 2}$, where $\mathbf{C}$ is the number of anomaly categories present in the dataset. Each row of this matrix corresponds to the ‘normal’ and ‘abnormal’ scores for a given category. For each category $c$, we denote the corresponding logit scores for the “normal” and “abnormal” classes as:
$\mathbf{z}_{c}=\begin{bmatrix}z_{c}^{n}\ z_{c}^{a}\end{bmatrix}\;\in\;\mathbb{R}^{2}.$ (5)
For each category $c$, we apply a normal vs. anomaly softmax per category to obtain the anomaly probability $p_{c}^{a}$. This type of separation into anomaly and normal cases is motivated by the observation that normal samples encompass a highly diverse range of conditions, making it difficult for the model to learn a coherent representation if they are treated as a single additional anomaly category. Unlike earlier works that risked treating ”normal” as just another anomaly type, we define the normal state by the absence of strong anomaly-specific characteristics, thereby avoiding distortion of learned representations. This is precisely why we designed the model to output the logits in this manner.
The category-wise anomaly probabilities serve as modulation signals that guide how the input representation is refined, enabling the model to softly encode the relevance of different anomaly categories based on their activation levels. Instead of selecting only the most likely category, as is typical in standard classification, the model derives a set of importance weights from the full anomaly probability distribution, reflecting its confidence across all categories.
Formally, we define the weights of each category as:
$w_{c}=\begin{cases}\dfrac{1}{C},&\text{if }S=\emptyset,\\[11.99998pt] \dfrac{\exp\bigl(z_{c}^{a}\bigr)}{\displaystyle\sum_{k=1}^{C}\exp\bigl(z_{k}^{a}\bigr)},&\text{if }S\neq\emptyset.\end{cases}$ (6)
where $S=\{\,c\mid p_{c}^{a}\geq\tau\,\}$ is the set of categories whose anomaly probability exceeds a threshold $\tau$. This allows the model to naturally separate normal inputs from anomaly categories without introducing a dedicated ”normal” prototype, enabling more faithful and effective learning that better reflects how anomalies are identified in real-world settings.
Category Prototype Injection. These weights are then used to softly combine the corresponding category embeddings, which act as prototypes representing each category, resulting in a feature representation that captures both shared and distinctive anomaly characteristics. Prototypes are learnable and arranged row-wise in the matrix:
$\mathbf{E}=\begin{bmatrix}\mathbf{e}_{1}^{\mathsf{T}}\\ \vdots\\ \mathbf{e}_{C}^{\mathsf{T}}\end{bmatrix}\;\in\;\mathbb{R}^{C\times d_{\mathrm{emb}}}.$ (7)
To encode contextual category semantics, we compute a soft category embedding vector $\mathbf{v}$ as a weighted sum over learnable category embeddings:
$\mathbf{v}=\sum_{c=1}^{C}w_{c}\,\mathbf{e}_{c},$ (8)
where $w_{c}$ represents the model’s estimated anomaly probability for the $c$-th category, and $\mathbf{e}_{c}$ is its corresponding embedding. This formulation allows the model to softly blend semantic cues across categories, capturing both confident and ambiguous class-related anomaly traits. During training, we enhance the soft category embedding $\mathbf{v}$ by adding the ground-truth class embedding, resulting in $\mathbf{v}_{\text{train}}=\mathbf{v}+\mathbf{e}_{y}$. This encourages class-aware specialization in the learned representations.
The embedding $\mathbf{v}$ is incorporated into the temporal-contextual features via a multi-head cross-attention mechanism. Specifically, $\mathbf{v}$ serves as the global query, while the temporal-contextual features act as both keys and values. This operation refines each segment feature $\mathbf{x}_{t}^{\text{tc}}$ into a category-aware refine feature $\mathbf{x}_{t}^{\text{ca}}$:
$\mathbf{x}_{t}^{\text{ca}}=\text{CrossAttn}(\mathbf{v},\mathbf{x}_{t}^{\text{tc}},\mathbf{x}_{t}^{\text{tc}}),$ (9)
effectively aligning low-level feature patterns with high-level category cues.
Finally, anomaly scoring is performed on the aligned features described above. This scoring part shares parameters with the earlier rough scoring function that was used for pooling features to be passed into the category classifier, enabling some degree of mutual learning during training.
We combine a Top‑$k$ MIL ranking loss (sultani2018real) $\mathcal{L}_{\mathrm{MIL}}$ to promote truly anomalous segments over normal ones, and an improved GMM‑based smoothing loss (wang2025learning) $\mathcal{L}_{\mathrm{GMM}}$. The original method fitted a Gaussian mixture by computing per‑video mean and variance across all segments; we enhance it by injecting our weighted category embedding into the mixture to more clearly reflect category‑specific traits. We also use a category classification loss $\mathcal{L}_{\mathrm{cat}}$, where Eq. 5 outputs a two-logit vector $z_{c}$ for each category $c$. For a video with ground-truth category $c^{*}$, we assign the target $t_{c}=[0,1]$ if $c=c^{*}$ and $t_{c}=[1,0]$ otherwise, and define $\mathcal{L}_{\mathrm{cat}}=\sum_{c=1}^{C}\mathrm{BCE}_{2}\!\left(\sigma(z_{c}),t_{c};w_{c}\right)$, with optional class weights $w_{c}$ to mitigate imbalance. This encourages the semantic embeddings to remain category-discriminative while retaining anomaly relevance.
The overall training objective is $\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{MIL}}+\lambda_{1}\,\mathcal{L}_{\mathrm{GMM}}+\lambda_{2}\,\mathcal{L}_{\mathrm{cat}}$.
Experiments
Experimental Setup
We evaluate our method on two widely used benchmark datasets for WSVAD: UCF-Crime(sultani2018real) and XD-Violence(wu2020not). UCF-Crime. For evaluation, we follow standard protocols. On UCF-Crime, we use the Area Under the Curve (AUC) and AUC for anomaly videos (AnoAUC) at the frame level. On XD-Violence, we use frame-level Average Precision (AP) and report mean Average Precision (mAP) across IoU thresholds from 0.1 to 0.5, along with the average score (AVG) for segment-level performance.
Approach Method Source
UCF-
AUC(%)
XD-
AP(%)
semi
-sup.
SVM baseline - 50.10 50.80
OCSVM (scholkopf1999support) NeurIPS 63.20 28.63
Hansan et al. (hasan2016learning) CVPR 51.20 31.25
weakly
-sup.
Ju et al. (ju2022prompting) ECCV 84.72 76.57
Sultani et al. (sultani2018real) CVPR 84.14 75.18
Wu et al. (wu2020not) ECCV 84.57 80.00
AVVD (wu2022weakly) TMM 82.45 78.10
RTFM (tian2021weakly) ICCV 85.66 78.27
DMU (zhou2023dual) AAAI 86.75 81.66
UMIL (lv2023unbiased) CVPR 86.75 -
CLIP-TSA (joo2023clip) ICIP 87.58 82.17
HSN (majhi2024human) CVIU 85.45 -
IFS-VAD (zhong2024inter) TCSVT 85.47 83.14
VadCLIP (wu2024vadclip) AAAI 88.02 84.57
PEMIL (chen2024prompt) CVPR 86.83 88.21
ReFLIP (dev2024reflip) TCSVT 88.57 85.81
CMHKF (wang2025cmhkf) ACL - 86.57
Ex-VAD (huangex) ICML 88.29 86.52
$\pi$-VAD (majhi2025just) CVPR 90.33 85.37
RefineVAD (Ours) 88.92 88.66
Table 1: Comparison of state-of-the-art methods on UCF-Crime and XD-Violence datasets. Best result is bolded and second best result is underlined.
Method mAP@IOU (%)
0.1 0.2 0.3 0.4 0.5 AVG
Random Baseline 0.21 0.14 0.04 0.02 0.01 0.08
Sultani et al. (sultani2018real) 5.73 4.41 2.69 1.93 1.44 3.24
AVVD (wu2022weakly) 10.27 7.01 6.25 3.42 3.29 6.05
VadCLIP (wu2024vadclip) 11.72 7.83 6.40 4.53 2.93 6.68
ITC (liu2024injecting) 13.54 9.24 7.45 5.46 3.79 7.90
ReFLIP (dev2024reflip) 14.23 10.34 9.32 7.54 6.81 9.62
Ex-VAD (huangex) 16.51 12.35 9.41 7.82 4.65 10.15
RefineVAD (Ours) 20.90 13.17 8.14 4.41 3.03 9.93
Table 2: Fine-grained comparisons on UCF-Crime.
MoTAR CORE AUC(%)
Category-Injection Soft-Classfication
84.60
85.43
87.28
87.85
88.89
Table 3: Ablation studies of our module on the UCF-Crime.
Implementation Details
We divide each video into 32 snippets and use pre-trained CLIP(VIT-L/14) (joo2023clip) as both the image and text encoder. For textual representation, InternVideo2.5 (wang2025internvideo25empoweringvideomllms) is utilized. The encoders are kept frozen during training. For the XD-Violence dataset, we train the model for a maximum of 30 epochs using the AdamW optimizer (LR: 3e-3, Batch Size: 64), where each mini-batch consisted of 32 normal and 32 abnormal samples. The loss weights were set to $\lambda_{1}=0.1$ and $\lambda_{2}=0.2$. The implementation used PyTorch on a single NVIDIA A5000 GPU.
Comparison on WVAD Benchmarks
We compare our performance with current state-of-the-art (SOTA) methods on UCF-Crime and XD-Violence datasets. According to Table 1, our model achieves 88.92% AUC on UCF-Crime and surpasses all previous weakly supervised learning models on XD-Violence with 88.66% AP. Regarding the segment-level evaluation in Table 2, our method achieves competitive mAP across IoU thresholds on UCF-Crime, ranking among the top existing approaches. Unlike prior works that explicitly optimize snippet-level classification, our framework predicts only a single video-level category, which is then used to guide temporal localization. While this design improves category-level agreement, it places our model at a structural disadvantage under higher IoU thresholds, where the metric emphasizes frame-accurate boundary localization rather than alignment with video-level predictions. The slight performance gap at high IoU thus arises from this evaluation–objective mismatch rather than from the capability of our approach. On XD-Violence, our method remains competitive and surpasses previous weakly supervised methods by 0.19%, validating the benefit of adaptive category-aware feature refinement via MoTAR and CORE.
Ablation Studies
To assess the contribution of each component, we conduct an ablation study on UCF-Crime (Table 3). Starting from the Base model (MLP with MIL), we sequentially add MoTAR, Category-Injection, and Soft Classification.
Introducing MoTAR provides a modest AUC gain (from 84.60% to 85.43%), showing the utility of motion-aware temporal adjustment. Notably, Category-Injection brings a much larger improvement, raising the AUC to 87.28% and delivering the most significant single-step boost. Finally, adding Soft Classification—where embeddings are formed not from a single top category but from a weighted sum over category probabilities—further refines the representation. Combining all components yields the best performance at 88.89%, confirming that each module is effective and that their integration provides complementary benefits for weakly supervised video anomaly detection.
Cross-Dataset Semantic Transfer Evaluation
To evaluate the transferability of the learned category semantics, we perform cross-dataset experiments on XD-Violence, as summarized in Table 4. Full training on XD-Violence serves as the upper-bound reference, achieving 88.66% AP. When freezing the CORE module (category classifier and semantic embeddings learned solely from UCF-Crime) and training the remaining modules, the model still attains 87.52% AP, indicating the UCF-Crime semantic space effectively encodes transferable category-level structure despite dataset differences. Finally, a strict zero-shot test—evaluating a UCF-Crime trained model directly on XD-Violence—achieves a robust 77.56% AP, demonstrating that the prototype-based semantics retain discriminative power across domains with mismatched category taxonomies and distinct visual distributions.
Collectively, these results show that our semantic representations extend beyond dataset-specific labels. Unlike prior category-driven approaches that required collecting new annotations for each target dataset and often degraded under mismatch, our compositional semantic space is label-efficient, scalable, and reusable without additional supervision, enabling practical cross-dataset deployment in weakly supervised video anomaly detection.
Method Category GT AP (%)
Full Training
88.66
Freezing CORE Module
(Category Classifier & Embeddings: Trained only on UCF)
87.52
Direct Cross-Domain Transfer
(Train: UCF-Crime, Test: XD-Violence, No Fine-Tuning)
77.56
Table 4: Semantic transfer of UCF-Crime category classifier and embedding on XD-Violence.
Figure 4: T-sne visualization of video features.
Figure 5: Example frame (top) and corresponding predicted scores with ground truth (bottom). Blue box indicates a normal frame, and red circles highlight anomalous regions.
Category Logits Visualization
To examine whether the predicted logit vectors are effectively separated according to semantic categories, we visualize the outputs of our soft category classifier using t-SNE. The logit vector corresponds to the category-wise activation output from the classifier, and similar vectors tend to be positioned close together in the embedding space.
As shown in the Figure 4, samples with similar semantic characteristics form distinct clusters. For instance, Arson, Explosion, and RoadAccidents are grouped together, where the classifier tends to focus on sudden background changes or scene-level phenomena. In contrast, Arrest, Assault, and Fighting are clustered with a focus on the interactions among multiple people. On the other hand, Shoplifting and Robbery form a third group, where the model mainly attends to the actions of a single individual.
This observation suggests that different anomaly categories require attention to different aspects of the scene (e.g., background changes, group activity, or individual behavior), with similar categories exhibiting similar logit activation distributions. These results validate the effectiveness of category-aware anomaly detection, where adapting the decision-making process based on the anomaly type leads to more accurate detection.
Conclusion
In this work, we proposed RefineVAD, a weakly-supervised framework that integrates motion-aware temporal recalibration and category-conditioned feature refinement to enhance anomaly localization. By jointly modeling motion patterns and semantic categories, RefineVAD captures both generic and category-specific anomaly cues, enabling more precise detection. Our soft classification mechanism guides the model toward meaningful temporal and semantic evidence, leading to improved performance across benchmarks. This approach also lays a foundation for future extensions toward fine-grained anomaly taxonomy, adaptive prototype refinement, and label-efficient cross-domain generalization.
Acknowledgments
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government(MSIT)(RS-2024-00456589) and Korea Planning & Evaluation Institute of Industrial Technology(KEIT) grant funded by the Korea government(MOTIE) (No. RS-2024-00442120, Development of AI technology capable of robustly recognizing abnormal and dangerous situations and behaviors during night and bad weather conditions)
RefineVAD: Semantic-Guided Feature Recalibration
for Weakly Supervised Video Anomaly Detection
Supplementary Material
Anonymous submission
Overview
This supplementary material provides additional technical details and supporting results that complement the main manuscript. It is organized as follows:
  • Section A Visualization of Category-wise Logit Activations Revealing Semantic Structure
  • Section B Performance Analysis of Category Classifier
  • Section C Category-aware Feature Injection Methods
  • Section D Visual Plots of Various Events
  • Section E Hyperparameters and Model Performance
  • Section F FLOPs Estimation and Computational Analysis of MoTAR
A. Visualization of Category-wise Logit Activations Revealing Semantic Structure
To complement the analysis presented in Section 4.3, we further illustrate the behavior of the category-aware logit outputs in Figure 6. This visualization highlights how the model not only correctly emphasizes the ground-truth category but also assigns relatively high activation scores to semantically related categories—a pattern consistently observed across multiple test instances.
Recall from Equation (5) that for each category $c$, the classifier outputs a two-dimensional logit vector:
$\mathbf{z}_{c}=\begin{bmatrix}z_{c}^{n}&z_{c}^{a}\end{bmatrix}\in\mathbb{R}^{2},$
where $z_{c}^{a}$ denotes the anomaly logit score associated with category $c$. These values form the basis of the category-wise anomaly confidence that guides the model’s decision process.
In Figure 6, we present per-category anomaly logit scores ($z_{c}^{a}$) as bar graphs for selected examples. The bars are color-coded as follows:
  • Red bars indicate the ground-truth category.
  • Purple bars denote categories within the same semantic group as the ground truth.
  • Blue bars represent all other unrelated categories.
This representation enables intuitive interpretation of how the model distributes its focus across semantically meaningful clusters. A particularly noteworthy and recurrent observation is that categories semantically close to the ground truth (e.g., Assault and Fighting, or Arson and Explosion) tend to receive significantly higher $z_{c}^{a}$ values than unrelated categories, even when they are not the correct label.
This phenomenon reinforces our earlier finding in Section 4.3, where t-SNE projections revealed that semantically similar categories naturally form tight clusters in the classifier’s embedding space. There, we observed that categories such as Arson, Explosion, and RoadAccidents gravitate toward one another based on background-centric cues, while categories like Assault, Arrest, and Fighting are grouped based on multi-person interactions.
For clarity, numerical labels are omitted in the figure for categories whose logit values fall below 1% of the maximum value. This ensures that the emphasis remains on the most relevant activations without visual clutter.
Overall, the bar-graph visualization in Figure 6 provides additional evidence that the model captures semantically structured inter-category relationships. These activation patterns further support the effectiveness of our category-aware anomaly detection framework.
(a) RoadAccidents follows Explosion with similarly abrupt background changes.
(b) RoadAccidents follows Explosion with similarly abrupt background changes.
(c) Assault and Arrest follow Fighting, reflecting intense interactions between individuals.
(d) Arson and Explosion follow RoadAccidents with similarly abrupt background changes
(e) Arson and Explosion follow RoadAccidents with similarly abrupt background changes.
(f) Shoplifting follows Robbery, sharing the characteristic of covert actions by a single person.
(g) Robbery follows Shoplifting, sharing the characteristic of covert actions by a single person.
(h) Shoplifting and Robbery show similarly high values, reflecting covert individual actions.
Figure 6: Visualization of inter-category similarities, grouped by shared visual dynamics and interaction types. Categories within each group exhibit closely aligned scores and patterns, indicating latent semantic consistency.
B. Performance Analysis of Category Classifier
In this section, we evaluate the standalone performance of the category classifier used in our proposed method, RefineVAD, on the UCF-Crime dataset. As discussed in the main paper, our model jointly trains a category classification branch and an anomaly detection branch via an interactive learning framework. One potential concern with this setup is that the category classifier might suffer performance degradation due to shared optimization with the anomaly scoring module.
To investigate this, we isolated the category classifier after training and tested it on a separate test set that includes only unseen videos, i.e., data that was not used during training. This allows us to assess the classifier’s generalization capability under joint training conditions.
Experimental Setup
  • Dataset: UCF-Crime
  • Categories: 13 anomalous and normal events
  • Evaluation: Classification accuracy on a held-out test set
  • Classifier Type: Shallow architecture, optimized for speed and integration with the main model
Quantitative Results
We compare the performance of the category classifier in two scenarios:
  1. 1.
    Joint Training (Ours): The classifier is trained alongside the anomaly detection module using combined losses.
  2. 2.
    Classification-Only Training: The classifier is trained independently using only category loss.
Table 5: Detailed Accuracy of the Category Classifier
Normal Accuracy Anomaly Accuracy
133 / 150 = 0.8867 68 / 140 = 0.4857
Interpretation and Discussion
Despite the simplicity of the classifier and the lack of exclusive optimization for category classification, our model achieves reasonable performance, especially for the ”Normal” class. An anomaly classification accuracy of 48.57% may appear low at first glance, but several factors justify this result:
  • Anomalous categories in UCF-Crime often contain overlapping visual semantics (e.g., Assault and Robbery), making fine-grained classification inherently ambiguous.
  • Our classifier was not trained in isolation. The shared learning process emphasizes anomaly detection, yet still preserves notable classification ability.
  • Even when trained using only the category loss, the classifier does not show a substantial difference in anomaly detection performance. This suggests that the difficulty arises from the inherent complexity of the task itself, rather than from the joint training strategy.
These results indicate that RefineVAD successfully integrates category supervision into the anomaly detection framework without introducing significant performance degradation. The proposed interactive learning approach preserves both interpretability and class-awareness while maintaining robust anomaly detection capability.
C. Category-aware Feature Injection Methods
Motivation
Although our model is trained under weak supervision (i.e., anomaly/normal binary labels), real-world anomaly videos contain diverse semantic categories such as Explosion, Fighting, or Arrest. While these categories are not directly supervised in the anomaly scoring process, they serve as informative metadata.
To exploit this latent information, we propose to inject category embeddings into snippet features. This enables the model to learn category-sensitive anomaly characteristics. We experimented with the following three strategies:
  1. 1.
    Linear Transformation (simple additive refinement)
  2. 2.
    Add & Norm with Feedforward Network (transformer-style residual enhancement)
  3. 3.
    Cross Attention Injection (ours) (context-aware dynamic alignment)
Formulations
Let $\mathbf{X}\in\mathbb{R}^{B\times T\times D}$ denote the snippet-level feature sequence, and let $\mathbf{e}_{c}\in\mathbb{R}^{d_{emb}}$ be the category embedding derived from predicted category logits.
1. Linear Transformation
A straightforward approach where the category embedding is linearly projected into the feature space and added to each snippet:
$\mathbf{F}_{\text{final}}=\mathbf{X}+\text{Linear}(\mathbf{e}_{c})$ (10)
Effect: Allows every snippet to share the same category signal; expected to provide coarse-level bias without temporal dynamics.
2. Add & Norm with FFN
Inspired by Transformer architecture, the category embedding is projected and added to snippet features, normalized, and enhanced via a feedforward layer:
$\mathbf{I}$ $=\text{LayerNorm}(\mathbf{X}+\mathbf{e}_{c})$ (11)
$\mathbf{F}_{\text{final}}$ $=\mathbf{I}+\text{FFN}(\mathbf{I})$ (12)
Effect: Introduces non-linearity and residual enhancement; expected to offer slightly more flexible feature refinement per time step.
3. Cross Attention (ours)
We treat the category embedding as a query and let it attend to the feature sequence via multi-head attention:
$\mathbf{C}$ $=\text{MultiHeadAttn}(\text{Query}=\mathbf{e}_{c},\ \text{Key}=\mathbf{X},\ \text{Value}=\mathbf{X})$ (13)
$\mathbf{F}_{\text{final}}$ $=\mathbf{X}+\mathbf{C}$ (14)
Effect: Provides dynamic, time-sensitive conditioning of snippet features based on category semantics; expected to learn nuanced, category-specific anomaly patterns.
Results and Comparison
We compare the three methods on two benchmark datasets. For UCF-Crime, we report Area Under Curve (AUC); for XD-Violence, we report Average Precision (AP).
Table 6: Comparison of category-aware injection methods. Metrics: AUC for UCF-Crime, AP for XD-Violence.
Injection Method UCF-Crime (AUC %) XD-Violence (AP %)
Linear Transformation 87.22 87.92
Add & Norm + FFN 87.95 86.83
Cross Attention (ours) 88.92 88.66
Our experiments demonstrate that while simple projection-based injection (Linear) or residual-style enhancement (Add & Norm + FFN) provide moderate gains, the cross-attention mechanism leads to the most substantial improvement. This confirms the importance of learning fine-grained interactions between category context and temporal dynamics in anomaly detection.
D. Visual Plots of Various Events
To further validate the effectiveness and robustness of our proposed method, we provide additional visualizations of frame-level anomaly scores in this appendix. These samples are supplementary to the limited number of visualizations presented in the main paper due to space constraints.
Each plot illustrates the frame-wise anomaly scores predicted by our model for a given video clip. The ground truth anomalous segments are highlighted with a red background to facilitate visual comparison. These visualizations demonstrate that our method consistently assigns higher anomaly scores within the ground truth anomalous regions, aligning well with temporal boundaries of anomalous events.
Coverage Across Categories
To ensure comprehensive evaluation, we include representative samples from all major anomaly categories in the UCF-Crime dataset. This wide coverage highlights the generalization ability of our model across diverse event types, and confirms that performance is not biased toward specific categories. For each category, at least one representative anomalous sample is presented.
(a) Arrest001
(b) Arson041
(c) Assault006
(d) Burglary024
(e) Burglary033
(f) Burglary037
(g) RoadAccidents012
(h) RoadAccidents016
(i) Shooting022
(j) Shoplifting005
(k) Stealing058
(l) Vandalism028
Figure 7: Visual examples from various abnormal video classes (4x3 layout)
Normal Case Analysis
In addition to anomalous cases, we include one visualization for a normal video sample. As expected, the anomaly scores for this sample remain consistently low across all frames. This example serves as a qualitative confirmation that our model does not falsely detect anomalies in normal videos, reinforcing the reliability of the scoring mechanism.
Figure 8: Example from normal video class
Purpose of Visualizations
These additional visualizations serve the following purposes:
  • To qualitatively demonstrate that our anomaly scoring mechanism correctly identifies and localizes anomalous segments across various event types.
  • To show that the model maintains low anomaly scores for normal videos, reducing the risk of false positives.
  • To provide transparent insight into how anomaly scores evolve over time, reinforcing the interpretability of the proposed method.
These visual results complement our quantitative evaluations and offer further evidence that our method robustly distinguishes anomalous behavior across diverse categories while suppressing spurious detections in normal scenes.
E. Hyperparameters and Model Performance
We conducted extensive hyperparameter optimization using a dedicated tuning library. For reproducibility, the random seed was fixed to 42 across all experiments. Approximately 1,000 trials were executed, using our proposed model RefineVAD as the backbone. The results for the XD-Violence datasets are summarized below.
Top Hyperparameter Importances:
The hyperparameter importances were estimated using a random forest regression model trained to predict the validation AP from the trial configurations. Each hyperparameter was treated as an input feature, and its relative contribution was measured based on the feature importance scores from the trained model.
Hyperparameter Importance
$d_{proj}$ 0.2873
$d_{emb}$ 0.2200
$\lambda_{cat}$ 0.1284
Learning Rate 0.1254
Epoch 0.0951
$\lambda_{reg}$ 0.0786
$\lambda_{GMM}$ 0.0287
Table 7: XD-Violence: Hyperparameter importance (based on AP)
Detailed Analysis of Key Hyperparameters:
We further investigated the three most influential hyperparameters — $d_{proj}$, $d_{emb}$, and $\lambda_{cat}$ — which exhibited the highest importance scores.
  • $d_{emb}$ denotes the dimensionality of the category embedding prototypes. This value determines how rich the semantic representation of each category can be, enabling the model to better distinguish between different types of normal and abnormal events.
  • $d_{proj}$ refers to the projection dimensionality used inside the CORE module when aligning features across different branches. It helps inject category-aware semantics into shared feature spaces.
  • $\lambda_{cat}$ controls the weight of the category classification loss term in the overall loss function. This term is essential in balancing the dual objectives of anomaly detection and category classification.
Figure 9: Maximum AP performance according to each hyperparameter. (Left) $d_{emb}$, (Middle) $d_{proj}$, (Right) $\lambda_{cat}$.
Both $d_{emb}$ and $d_{proj}$ had a considerable impact on performance. Since they are directly involved in injecting category-aware semantic knowledge into feature representations, they influence how effectively the model can reason over category-conditioned patterns.
In contrast, $\lambda_{cat}$ affects the trade-off between anomaly detection and category classification. Thus, we conducted a finer-grained search over a narrower range to identify the optimal balance between these tasks. This highlights the importance of careful tuning in multi-objective setups.
F. Computational Analysis of MoTAR
We report the the computational efficiency of the proposed MoTAR module, including parameter count, floating point operations (FLOPs), memory usage, and inference speed. FLOPs represent the number of scalar floating point operations required during a forward pass and are commonly used as a hardware-agnostic measure of computational cost. In our calculation, we follow standard conventions used in prior work: a matrix multiplication between tensors of shape $M\times K]$ and $K\times N]$ is assumed to require $2\cdot M\cdot K\cdot N$ FLOPs, and simple element-wise operations such as activation functions and additions are counted as 1 FLOP per element.
Temporal Module Params FLOPs Memory(GB) Speed(ms)
MTN (tian2021weakly) 14.17M 1.2G 0.2419 1.500
LGTAdapter (wu2024vadclip) 18.63M 13.18G 0.2985 2.143
MoTAR (Ours) 11.03M 242M 0.1905 0.856
Table 8: Computational efficiency comparison across temporal modules.
The MoTAR module consists of three main components: (1) a lightweight MLP-based ShiftPredictor, (2) an adaptive temporal shifting mechanism, and (3) a multi-layer TemporalTransformer block. The ShiftPredictor computes variance over frame differences and passes the result through a 3-layer MLP (768→32→16→1) with GELU activations. For a typical input of shape [B, T, D] = [64, 32, 768], the entire MoTAR module contributes approximately 11.03M Params, 242M FLOPs, 0.1905GB Memory and 0.856ms inference latency including both the variance computation and MLP forward pass.
The adaptive shifting operation is input-dependent and implemented through per-timestep control-flow logic involving conditionals and dynamic indexing. Since this operation cannot be statically analyzed and varies at runtime, we follow prior work in excluding such components from FLOPs estimation. Importantly, unlike conventional temporal modeling approaches such as Graph Convolution, MoTAR’s shifting mechanism does not involve learnable parameters or floating-point operations during inference, contributing to significantly lower computational cost. This non-parametric and zero-computation design is crucial to MoTAR’s overall resource efficiency, directly contributing to the significantly reduced parameter count and minimal inference overhead.
The remaining computational cost mainly arises from the Temporal Transformer, which consists of two stacked TransformerEncoderLayers, each comprising a multi-head self-attention (MHSA) block and a feed-forward network. For a hidden dimension D = 768, sequence length T = 64, and number of heads H = 4, the total FLOPs for the transformer component amounts to approximately 121M FLOPs. The vast majority of the 11.03M parameters are contributed by the weight matrices within the Temporal Transformer’s MHSA and FFN blocks. Correspondingly, the 0.1905GB memory usage is dominated by storing the module’s weights and the intermediate activation tensors generated during the forward pass.
Summing up the FLOPs from the ShiftPredictor and TemporalTransformer, yields a total of about 242M FLOPs per forward pass (excluding the adaptive shift logic). This analysis helps isolate the dominant computational contributors within the module and guides future design trade-offs between adaptivity and efficiency. As shown in Tab. 8, MoTAR achieves substantial resource efficiency across all measured metrics. The combined effect of low FLOPs, the optimized parameter count (11.03M), and minimal intermediate tensor storage results in a low memory footprint (0.1905GB), which ultimately translates directly into the fastest observed inference speed (0.856ms). This confirms that MoTAR successfully achieves its adaptive functionality with minimal overhead.

### v2

@@
 On XD-Violence, we use frame-level Average Precision (AP) and report mean Average Precision (mAP) across IoU thresholds from 0.1 to 0.5, along with the average score (AVG) for segment-level performance.
-Approach Method Source
-UCF-
-AUC(%)
-XD-
-AP(%)
-semi
--sup.
-SVM baseline - 50.10 50.80
+Approach Method Source UCF- AUC(%) XD- AP(%)
+semi -sup. SVM baseline - 50.10 50.80
 OCSVM (scholkopf1999support) NeurIPS 63.20 28.63
 Hansan et al. (hasan2016learning) CVPR 51.20 31.25
-weakly
--sup.
-Ju et al. (ju2022prompting) ECCV 84.72 76.57
+weakly -sup. Ju et al. (ju2022prompting) ECCV 84.72 76.57
 Sultani et al. (sultani2018real) CVPR 84.14 75.18
@@
 Method Category GT AP (%)
-Full Training
-88.66
-Freezing CORE Module
-(Category Classifier & Embeddings: Trained only on UCF)
-87.52
-Direct Cross-Domain Transfer
-(Train: UCF-Crime, Test: XD-Violence, No Fine-Tuning)
-77.56
+Full Training 88.66
+Freezing CORE Module (Category Classifier & Embeddings: Trained only on UCF) 87.52
+Direct Cross-Domain Transfer (Train: UCF-Crime, Test: XD-Violence, No Fine-Tuning) 77.56
 Table 4: Semantic transfer of UCF-Crime category classifier and embedding on XD-Violence.
-Figure 4: T-sne visualization of video features.
-Figure 5: Example frame (top) and corresponding predicted scores with ground truth (bottom). Blue box indicates a normal frame, and red circles highlight anomalous regions.
+Figure 4: t-SNE visualization of logit features, where semantically similar categories form meaningful clusters.
+Figure 5: Example frame (top) and corresponding predicted scores with ground truth (bottom).
+The blue curve represents the predicted scores, and the red shaded regions indicate the ground-truth anomalous intervals.
 Category Logits Visualization
@@
 Our soft classification mechanism guides the model toward meaningful temporal and semantic evidence, leading to improved performance across benchmarks.
+In addition, the resulting category-aware feature space provides improved interpretability and demonstrates strong transferability across datasets.
 This approach also lays a foundation for future extensions toward fine-grained anomaly taxonomy, adaptive prototype refinement, and label-efficient cross-domain generalization.
@@
 RS-2024-00442120, Development of AI technology capable of robustly recognizing abnormal and dangerous situations and behaviors during night and bad weather conditions)
-RefineVAD: Semantic-Guided Feature Recalibration
-for Weakly Supervised Video Anomaly Detection
-Supplementary Material
-Anonymous submission
-Overview
-This supplementary material provides additional technical details and supporting results that complement the main manuscript. It is organized as follows:
-  • Section A Visualization of Category-wise Logit Activations Revealing Semantic Structure
-  • Section B Performance Analysis of Category Classifier
-  • Section C Category-aware Feature Injection Methods
-  • Section D Visual Plots of Various Events
-  • Section E Hyperparameters and Model Performance
-  • Section F FLOPs Estimation and Computational Analysis of MoTAR
-A. Visualization of Category-wise Logit Activations Revealing Semantic Structure
-To complement the analysis presented in Section 4.3, we further illustrate the behavior of the category-aware logit outputs in Figure 6.
-This visualization highlights how the model not only correctly emphasizes the ground-truth category but also assigns relatively high activation scores to semantically related categories—a pattern consistently observed across multiple test instances.
-Recall from Equation (5) that for each category $c$, the classifier outputs a two-dimensional logit vector:
-$\mathbf{z}_{c}=\begin{bmatrix}z_{c}^{n}&z_{c}^{a}\end{bmatrix}\in\mathbb{R}^{2},$
-where $z_{c}^{a}$ denotes the anomaly logit score associated with category $c$. These values form the basis of the category-wise anomaly confidence that guides the model’s decision process.
-In Figure 6, we present per-category anomaly logit scores ($z_{c}^{a}$) as bar graphs for selected examples. The bars are color-coded as follows:
-  • Red bars indicate the ground-truth category.
-  • Purple bars denote categories within the same semantic group as the ground truth.
-  • Blue bars represent all other unrelated categories.
-This representation enables intuitive interpretation of how the model distributes its focus across semantically meaningful clusters.
-A particularly noteworthy and recurrent observation is that categories semantically close to the ground truth (e.g., Assault and Fighting, or Arson and Explosion) tend to receive significantly higher $z_{c}^{a}$ values than unrelated categories, even when they are not the correct label.
-This phenomenon reinforces our earlier finding in Section 4.3, where t-SNE projections revealed that semantically similar categories naturally form tight clusters in the classifier’s embedding space.
-There, we observed that categories such as Arson, Explosion, and RoadAccidents gravitate toward one another based on background-centric cues, while categories like Assault, Arrest, and Fighting are grouped based on multi-person interactions.
-For clarity, numerical labels are omitted in the figure for categories whose logit values fall below 1% of the maximum value.
-This ensures that the emphasis remains on the most relevant activations without visual clutter.
-Overall, the bar-graph visualization in Figure 6 provides additional evidence that the model captures semantically structured inter-category relationships.
-These activation patterns further support the effectiveness of our category-aware anomaly detection framework.
-(a) RoadAccidents follows Explosion with similarly abrupt background changes.
-(b) RoadAccidents follows Explosion with similarly abrupt background changes.
-(c) Assault and Arrest follow Fighting, reflecting intense interactions between individuals.
-(d) Arson and Explosion follow RoadAccidents with similarly abrupt background changes
-(e) Arson and Explosion follow RoadAccidents with similarly abrupt background changes.
-(f) Shoplifting follows Robbery, sharing the characteristic of covert actions by a single person.
-(g) Robbery follows Shoplifting, sharing the characteristic of covert actions by a single person.
-(h) Shoplifting and Robbery show similarly high values, reflecting covert individual actions.
-Figure 6: Visualization of inter-category similarities, grouped by shared visual dynamics and interaction types.
-Categories within each group exhibit closely aligned scores and patterns, indicating latent semantic consistency.
-B. Performance Analysis of Category Classifier
-In this section, we evaluate the standalone performance of the category classifier used in our proposed method, RefineVAD, on the UCF-Crime dataset.
-As discussed in the main paper, our model jointly trains a category classification branch and an anomaly detection branch via an interactive learning framework.
-One potential concern with this setup is that the category classifier might suffer performance degradation due to shared optimization with the anomaly scoring module.
-To investigate this, we isolated the category classifier after training and tested it on a separate test set that includes only unseen videos, i.e., data that was not used during training.
-This allows us to assess the classifier’s generalization capability under joint training conditions.
-Experimental Setup
-  • Dataset: UCF-Crime
-  • Categories: 13 anomalous and normal events
-  • Evaluation: Classification accuracy on a held-out test set
-  • Classifier Type: Shallow architecture, optimized for speed and integration with the main model
-Quantitative Results
-We compare the performance of the category classifier in two scenarios:
-  1. 1.
-    Joint Training (Ours): The classifier is trained alongside the anomaly detection module using combined losses.
-  2. 2.
-    Classification-Only Training: The classifier is trained independently using only category loss.
-Table 5: Detailed Accuracy of the Category Classifier
-Normal Accuracy Anomaly Accuracy
-133 / 150 = 0.8867 68 / 140 = 0.4857
-Interpretation and Discussion
-Despite the simplicity of the classifier and the lack of exclusive optimization for category classification, our model achieves reasonable performance, especially for the ”Normal” class.
-An anomaly classification accuracy of 48.57% may appear low at first glance, but several factors justify this result:
-  • Anomalous categories in UCF-Crime often contain overlapping visual semantics (e.g., Assault and Robbery), making fine-grained classification inherently ambiguous.
-  • Our classifier was not trained in isolation. The shared learning process emphasizes anomaly detection, yet still preserves notable classification ability.
-  • Even when trained using only the category loss, the classifier does not show a substantial difference in anomaly detection performance.
-This suggests that the difficulty arises from the inherent complexity of the task itself, rather than from the joint training strategy.
-These results indicate that RefineVAD successfully integrates category supervision into the anomaly detection framework without introducing significant performance degradation.
-The proposed interactive learning approach preserves both interpretability and class-awareness while maintaining robust anomaly detection capability.
-C. Category-aware Feature Injection Methods
-Motivation
-Although our model is trained under weak supervision (i.e., anomaly/normal binary labels), real-world anomaly videos contain diverse semantic categories such as Explosion, Fighting, or Arrest.
-While these categories are not directly supervised in the anomaly scoring process, they serve as informative metadata.
-To exploit this latent information, we propose to inject category embeddings into snippet features.
-This enables the model to learn category-sensitive anomaly characteristics.
-We experimented with the following three strategies:
-  1. 1.
-    Linear Transformation (simple additive refinement)
-  2. 2.
-    Add & Norm with Feedforward Network (transformer-style residual enhancement)
-  3. 3.
-    Cross Attention Injection (ours) (context-aware dynamic alignment)
-Formulations
-Let $\mathbf{X}\in\mathbb{R}^{B\times T\times D}$ denote the snippet-level feature sequence, and let $\mathbf{e}_{c}\in\mathbb{R}^{d_{emb}}$ be the category embedding derived from predicted category logits.
-1. Linear Transformation
-A straightforward approach where the category embedding is linearly projected into the feature space and added to each snippet:
-$\mathbf{F}_{\text{final}}=\mathbf{X}+\text{Linear}(\mathbf{e}_{c})$ (10)
-Effect: Allows every snippet to share the same category signal; expected to provide coarse-level bias without temporal dynamics.
-2. Add & Norm with FFN
-Inspired by Transformer architecture, the category embedding is projected and added to snippet features, normalized, and enhanced via a feedforward layer:
-$\mathbf{I}$ $=\text{LayerNorm}(\mathbf{X}+\mathbf{e}_{c})$ (11)
-$\mathbf{F}_{\text{final}}$ $=\mathbf{I}+\text{FFN}(\mathbf{I})$ (12)
-Effect: Introduces non-linearity and residual enhancement; expected to offer slightly more flexible feature refinement per time step.
-3. Cross Attention (ours)
-We treat the category embedding as a query and let it attend to the feature sequence via multi-head attention:
-$\mathbf{C}$ $=\text{MultiHeadAttn}(\text{Query}=\mathbf{e}_{c},\ \text{Key}=\mathbf{X},\ \text{Value}=\mathbf{X})$ (13)
-$\mathbf{F}_{\text{final}}$ $=\mathbf{X}+\mathbf{C}$ (14)
-Effect: Provides dynamic, time-sensitive conditioning of snippet features based on category semantics; expected to learn nuanced, category-specific anomaly patterns.
-Results and Comparison
-We compare the three methods on two benchmark datasets. For UCF-Crime, we report Area Under Curve (AUC); for XD-Violence, we report Average Precision (AP).
-Table 6: Comparison of category-aware injection methods. Metrics: AUC for UCF-Crime, AP for XD-Violence.
-Injection Method UCF-Crime (AUC %) XD-Violence (AP %)
-Linear Transformation 87.22 87.92
-Add & Norm + FFN 87.95 86.83
-Cross Attention (ours) 88.92 88.66
-Our experiments demonstrate that while simple projection-based injection (Linear) or residual-style enhancement (Add & Norm + FFN) provide moderate gains, the cross-attention mechanism leads to the most substantial improvement.
-This confirms the importance of learning fine-grained interactions between category context and temporal dynamics in anomaly detection.
-D. Visual Plots of Various Events
-To further validate the effectiveness and robustness of our proposed method, we provide additional visualizations of frame-level anomaly scores in this appendix.
-These samples are supplementary to the limited number of visualizations presented in the main paper due to space constraints.
-Each plot illustrates the frame-wise anomaly scores predicted by our model for a given video clip.
-The ground truth anomalous segments are highlighted with a red background to facilitate visual comparison.
-These visualizations demonstrate that our method consistently assigns higher anomaly scores within the ground truth anomalous regions, aligning well with temporal boundaries of anomalous events.
-Coverage Across Categories
-To ensure comprehensive evaluation, we include representative samples from all major anomaly categories in the UCF-Crime dataset.
-This wide coverage highlights the generalization ability of our model across diverse event types, and confirms that performance is not biased toward specific categories.
-For each category, at least one representative anomalous sample is presented.
-(a) Arrest001
-(b) Arson041
-(c) Assault006
-(d) Burglary024
-(e) Burglary033
-(f) Burglary037
-(g) RoadAccidents012
-(h) RoadAccidents016
-(i) Shooting022
-(j) Shoplifting005
-(k) Stealing058
-(l) Vandalism028
-Figure 7: Visual examples from various abnormal video classes (4x3 layout)
-Normal Case Analysis
-In addition to anomalous cases, we include one visualization for a normal video sample.
-As expected, the anomaly scores for this sample remain consistently low across all frames.
-This example serves as a qualitative confirmation that our model does not falsely detect anomalies in normal videos, reinforcing the reliability of the scoring mechanism.
-Figure 8: Example from normal video class
-Purpose of Visualizations
-These additional visualizations serve the following purposes:
-  • To qualitatively demonstrate that our anomaly scoring mechanism correctly identifies and localizes anomalous segments across various event types.
-  • To show that the model maintains low anomaly scores for normal videos, reducing the risk of false positives.
-  • To provide transparent insight into how anomaly scores evolve over time, reinforcing the interpretability of the proposed method.
-These visual results complement our quantitative evaluations and offer further evidence that our method robustly distinguishes anomalous behavior across diverse categories while suppressing spurious detections in normal scenes.
-E. Hyperparameters and Model Performance
-We conducted extensive hyperparameter optimization using a dedicated tuning library.
-For reproducibility, the random seed was fixed to 42 across all experiments.
-Approximately 1,000 trials were executed, using our proposed model RefineVAD as the backbone.
-The results for the XD-Violence datasets are summarized below.
-Top Hyperparameter Importances:
-The hyperparameter importances were estimated using a random forest regression model trained to predict the validation AP from the trial configurations.
-Each hyperparameter was treated as an input feature, and its relative contribution was measured based on the feature importance scores from the trained model.
-Hyperparameter Importance
-$d_{proj}$ 0.2873
-$d_{emb}$ 0.2200
-$\lambda_{cat}$ 0.1284
-Learning Rate 0.1254
-Epoch 0.0951
-$\lambda_{reg}$ 0.0786
-$\lambda_{GMM}$ 0.0287
-Table 7: XD-Violence: Hyperparameter importance (based on AP)
-Detailed Analysis of Key Hyperparameters:
-We further investigated the three most influential hyperparameters — $d_{proj}$, $d_{emb}$, and $\lambda_{cat}$ — which exhibited the highest importance scores.
-  • $d_{emb}$ denotes the dimensionality of the category embedding prototypes.
-This value determines how rich the semantic representation of each category can be, enabling the model to better distinguish between different types of normal and abnormal events.
-  • $d_{proj}$ refers to the projection dimensionality used inside the CORE module when aligning features across different branches.
-It helps inject category-aware semantics into shared feature spaces.
-  • $\lambda_{cat}$ controls the weight of the category classification loss term in the overall loss function.
-This term is essential in balancing the dual objectives of anomaly detection and category classification.
-Figure 9: Maximum AP performance according to each hyperparameter. (Left) $d_{emb}$, (Middle) $d_{proj}$, (Right) $\lambda_{cat}$.
-Both $d_{emb}$ and $d_{proj}$ had a considerable impact on performance.
-Since they are directly involved in injecting category-aware semantic knowledge into feature representations, they influence how effectively the model can reason over category-conditioned patterns.
-In contrast, $\lambda_{cat}$ affects the trade-off between anomaly detection and category classification.
-Thus, we conducted a finer-grained search over a narrower range to identify the optimal balance between these tasks.
-This highlights the importance of careful tuning in multi-objective setups.
-F. Computational Analysis of MoTAR
-We report the the computational efficiency of the proposed MoTAR module, including parameter count, floating point operations (FLOPs), memory usage, and inference speed.
-FLOPs represent the number of scalar floating point operations required during a forward pass and are commonly used as a hardware-agnostic measure of computational cost.
-In our calculation, we follow standard conventions used in prior work: a matrix multiplication between tensors of shape $M\times K]$ and $K\times N]$ is assumed to require $2\cdot M\cdot K\cdot N$ FLOPs, and simple element-wise operations such as activation functions and additions are counted as 1 FLOP per element.
-Temporal Module Params FLOPs Memory(GB) Speed(ms)
-MTN (tian2021weakly) 14.17M 1.2G 0.2419 1.500
-LGTAdapter (wu2024vadclip) 18.63M 13.18G 0.2985 2.143
-MoTAR (Ours) 11.03M 242M 0.1905 0.856
-Table 8: Computational efficiency comparison across temporal modules.
-The MoTAR module consists of three main components: (1) a lightweight MLP-based ShiftPredictor, (2) an adaptive temporal shifting mechanism, and (3) a multi-layer TemporalTransformer block.
-The ShiftPredictor computes variance over frame differences and passes the result through a 3-layer MLP (768→32→16→1) with GELU activations.
-For a typical input of shape [B, T, D] = [64, 32, 768], the entire MoTAR module contributes approximately 11.03M Params, 242M FLOPs, 0.1905GB Memory and 0.856ms inference latency including both the variance computation and MLP forward pass.
-The adaptive shifting operation is input-dependent and implemented through per-timestep control-flow logic involving conditionals and dynamic indexing.
-Since this operation cannot be statically analyzed and varies at runtime, we follow prior work in excluding such components from FLOPs estimation.
-Importantly, unlike conventional temporal modeling approaches such as Graph Convolution, MoTAR’s shifting mechanism does not involve learnable parameters or floating-point operations during inference, contributing to significantly lower computational cost.
-This non-parametric and zero-computation design is crucial to MoTAR’s overall resource efficiency, directly contributing to the significantly reduced parameter count and minimal inference overhead.
-The remaining computational cost mainly arises from the Temporal Transformer, which consists of two stacked TransformerEncoderLayers, each comprising a multi-head self-attention (MHSA) block and a feed-forward network.
-For a hidden dimension D = 768, sequence length T = 64, and number of heads H = 4, the total FLOPs for the transformer component amounts to approximately 121M FLOPs.
-The vast majority of the 11.03M parameters are contributed by the weight matrices within the Temporal Transformer’s MHSA and FFN blocks.
-Correspondingly, the 0.1905GB memory usage is dominated by storing the module’s weights and the intermediate activation tensors generated during the forward pass.
-Summing up the FLOPs from the ShiftPredictor and TemporalTransformer, yields a total of about 242M FLOPs per forward pass (excluding the adaptive shift logic).
-This analysis helps isolate the dominant computational contributors within the module and guides future design trade-offs between adaptivity and efficiency.
-As shown in Tab.
-8, MoTAR achieves substantial resource efficiency across all measured metrics.
-The combined effect of low FLOPs, the optimized parameter count (11.03M), and minimal intermediate tensor storage results in a low memory footprint (0.1905GB), which ultimately translates directly into the fastest observed inference speed (0.856ms).
-This confirms that MoTAR successfully achieves its adaptive functionality with minimal overhead.